boy standing at a corner noticed a thief on bike moving with 54 km/hr and after some time a police jeep crosses him with speed 72 km/hr. if jeep catches the thief at a distance of 1 km from the boy then the time gap between bike & jeep will be when they cross the boy :-

Grade:12

1 Answers

Dhruv
6 years ago
Given,Speed of theif=54km/hrSpeed of police=72km/hrNow,If theif travel 54km in=1hrThen theif travel 1km in =1/54hr =66.66secNow,If police travel 72km in =1hrThen police travel 1km in = 1/72hr =50secTime difference between theif and police after crossing the boy =66.66sec-50sec =16.6sec